A ball is thrown horizontally from the top of
a building 120 m high. The ball strikes the
ground 63 m horizontally from the point of
release. What is the speed of the ball just before it
strikes the ground?
Answer in units of m/s.
1
Expert's answer
2019-10-07T10:47:18-0400

Let the height of the building be "H = 120m" and the horizontal travel distance "L = 63 m".

The equations of motion are as follows:

"v_x(t) = v_{0x}", "v_y(t) = v_{0y} - g t = - g t" (no initial vertical component of velocity),

"H = \frac{g t^2}{2}", "L = v_{0x} t" .

From the last two equations, "t =\sqrt{\frac{2 H}{g}}", "v_{0x} = \frac{L}{t} = \frac{L \sqrt{g}}{\sqrt{2 H}}". The speed before hitting the ground is "v = \sqrt{v_x^2 + v_y^2} = \sqrt{v_{0x}^2 + (g t)^2} = \sqrt{\frac{g L^2}{2 H} + 2 g H} \approx 50.2 \frac{m}{s}".